if the total bill is $1000 and the tax rate is 6.25% what was the original amount of the bill?
What is the formula used to solve this problem?

Found 2 solutions by nerdybill, stanbon:
Answer by nerdybill(7384)

if the total bill is $1000 and the tax rate is 6.25% what was the original amount of the bill?
What is the formula used to solve this problem?
Let x = original amount of bill
then
x + 6.25/100(x) = 1000
x + .0625(x) = 1000
x(1+.0625) = 1000
x(1.0625) = 1000
x = 1000/1.0625 (formula)
x = $941.18

Answer by stanbon(75887)

if the total bill is $1000 and the tax rate is 6.25% what was the original amount of the bill?
What is the formula used to solve this problem?
--------
original cost + tax = bill
x + 0.0625x = 1000
----
1.0625x = 1000
x = 1000/10625
x = $941.18
----------------
Cheers,
Stan H.
==================